\ The following data are accumulated by Eco Labs, Inc. in evaluating two competing capital investment...

\

  1. The following data are accumulated by Eco Labs, Inc. in evaluating two competing capital investment proposals:

    Testing Equipment Vehicle
    Amount of investment $60,000 $92,000
    Useful life 4 years 7 years
    Estimated residual value 0 0
    Estimated total income over the useful life $9,000 $32,200

    Determine the expected average rate of return for each proposal. If required, round your answers to one decimal place.

    Testing Equipment %
    Vehicle %
Answer #1

Average rate of return = Annual net income / Average investment

Testing equipment = (9000/4) / (60,000/2)

= 2250 / 30,000

= 7.5 %

Vehicle = (32,200/7) / (92,000/2)

= 4600 / 46,000

= 10%

Comment if you face any issues